Write an equation for the line that is perpendicular to the given line and that passes through a given point.
4x-12y=2;(10,-1)
Here is what I did:
-12y=-4x+2 (divided every thing by -12)
y= 3/4x+1/6 (negative reciprocal is -4/3)
y=mx+b
-1=(-4/3)*10+b
-1=-40+b
-1+40=b
39=b (plug in "b")
the equation is: y=-4/3x+39 (I am not sure if this is right)

Answer by nerdybill(7384)

-12y=-4x+2 (divided every thing by -12)
y= 3/4x+1/6 (negative reciprocal is -4/3)
y=mx+b
-1=(-4/3)*10+b
-1=-40+b <--WRONG, SHOULD BE:
-1=(-40/3)+b
-1+40/3=b
-3/3+40/3=b
37/3 = B (plug in "b")
the equation is: y=-4/3x+37/3
.
Or, you could have plugged it all into the "point-slope" form:
y - y1 = m(x - x1)
y - (-1) = (-4/3)(x - 10)
y + 1 = (-4/3)x + 40/3
y = (-4/3)x + 40/3 - 1
y = (-4/3)x + 40/3 - 3/3
y = (-4/3)x + 37/3